### v1

@@ V-A Color Correlations
$\rho(\mathbf{g})=\frac{\langle\delta(\mathbf{r},\mathbf{g}),\delta(\mathbf{g},\mathbf{b})\rangle}{\|\delta(\mathbf{r},\mathbf{g})\|\|\delta(\mathbf{g},\mathbf{b})\|}\enspace.$ (4)
@@
Fig. 2 (c) examines correlations induced by traditional JPEG compression. JPEG compression uses a similar preprocessing chain by converting RGB to YCbCr and applying chroma subsampling. The uncompressed patches are compressed with JPEG quality factors $5$, $20$, $50$, and $100$. During compression, we ensure that the same 4:2:0 chroma subsampling is used as for JPEG AI. For reasonably high quality factor of $100$, the color correlations slightly increase. With stronger compression of quality factor $50$ the color correlations are almost as much increased as it could be observed for the JPEG AI preprocessing. This is plausible, given the overall similarity of the preprocessing. It is interesting to note that for very strong compressions of quality factors $20$ or even only $5$, the correlations are weakened. We did not further investigate into this effect, but we hypothesize that the large quantization factors and associated large rounding errors act as a decorrelator between the color channels.
Fig. 2 (d) examines correlations in synthetic images. Similar correlations in synthetic images have been reported earlier for some color spaces [40]. Here, we show the correlations for synthetic images from the four diffusion-based image generators DALL-E2, Midjourney-V5 (Midj-V5), Firefly, and Stable Diffusion XL (SDXL). Images from DALL-E2 and SDXL exhibit remarkably low correlations in the range of $[-0.1,+0.15]$. In contrast, images from Midjourney-V5 and Firefly exhibit correlations that are comparable to traditional JPEG compression and JPEG AI compression at very high bitrates. It may be a coincidence that these generators produce images with more natural color appearance, while DALL-E2 and SDXL tend to generate images that oftentimes look oversaturated. In any case, none of the observed correlations from traditional JPEG or synthetic image generators is as large as the correlations for stronger JPEG AI compression.
@@
Figure 4: PSNR between compression $k-1$ and $k$ per bitrate. We analyze JPEG AI (left), the AI codec by Ballé et al. [7] (middle) and JPEG (right). See text for details.
Figure 5: Dependencies between rate $r(k)$ and PSNR $p_{\mathrm{inc}}(k)$ during $k$ compression runs to differentiate between single and recompressed JPEG AI images. In particular, $r(1)$ and the difference $r(3)-r(2)$ are more discriminative than the others.
@@ V-B Rate-Distortion Cue
The image and a first version of its associated rate-distortion diagrams are shown in Fig. 3. The left, middle, and right plots show the diagrams for JPEG AI, a further AI codec by Ballé et al. [7], denoted as Ballé et al and traditional JPEG. The $x$-axis indicates the bitrate of the compressed image. JPEG AI allows to directly specify a target bitrate, and for JPEG and Ballé et al. we varied the compression quality to achieve compressions with bitrates between $0.5$ and $3.5$. An image is compressed between $1$ (darkblue) to $4$ times (lightblue). The $y$-axis indicates the PSNR between the uncompressed input image and the image after $k$ compression steps. For traditional JPEG, repeated compressions barely lead to a change in PSNR, particularly at low bitrates. Only at a bitrate of $2.0$ or higher, PSNRs are decreased by iterated compression. For JPEG AI and Ballé et al., a loss in PSNR upon repeated compression is noticable across the whole range of investigated bitrates.
It is insightful to study a second type of rate-distortion diagram, which is shown in Fig. 4. Here, the $y$-axis indicates the PSNR between two consecutive compressions $k-1$ and $k$. Hence, PSNR is lowest between the original image and the 1st compression, because this compression step changes the image most. Subsequent recompressions have smaller impact, hence the PSNR between compression steps $k-1$ and $k$ is larger. This behavior is analogous for all three codecs, and has also been used in classical works on JPEG recompression forensics [26]. The PSNR differences of both AI codecs increase with increasing bitrates. In contrast, JPEG’s PSNRs are very large at low bitrates, indicating that recompression does not change the image much, but flatten for increasing bitrates somewhere between PSNRs of 45 to 60. As a sidenote, the differences in PSNR in Fig. 4 do not necessarily lead to a noticable difference in PSNR in Fig. 3 (probably best seen for low JPEG bitrates), which is rooted in the instable block phenomenon, where individual JPEG blocks keep changing upon recompression [26, 29].
@@
  • 4 features are obtained from PSNRs. Let $p_{\mathrm{inp}}(k)$ denote the PSNR between the input image and the $k$-th recompression (as in Fig. 3), and let $p_{\mathrm{inc}}(k)$ denote the incremental PSNR between the $k-1$-th and the $k$-th recompression (as in Fig. 4). Then, the features are $p_{\mathrm{inp}}(2)$, $p_{\mathrm{inp}}(3)$, $p_{\mathrm{inc}}(2)$, and $p_{\mathrm{inc}}(3)$.
  • 4 features from the differences $(r(3)-r(2))$, $(r(2)-r(1))$, $(p_{\mathrm{inp}}(3)-p_{\mathrm{inp}}(2))$, and $(p_{\mathrm{inc}}(3)-p_{\mathrm{inc}}(2))$.
@@
Figure 5 shows the dependencies between selected features during the compression runs. We observe that the calculated PSNR values and the rate of the three compression runs have a potential to differentiate between single and recompressed JPEG AI images. Notably, $r(1)$ and the difference $(r(3)-r(2))$ are more discriminative than $(r(2)-r(1))$ and $(p_{\mathrm{inc}}(3)-p_{\mathrm{inc}}(2))$.
@@ V-C Quantization Cue
These correlations capture two types of properties. First, they are sensitive to zero entries (as a zero entry in one vector cancels the entry in the other vector), and second they tend to reward similar step sizes. Zero entries are arguably most relevant for detecting lossy compression (aiming to maximize sequences of zeros for the arithmetic encoding), but also off-zero entries play an important role. To show this, we also compute a correlation on binarized vectors, i.e., we truncate all non-zero entries to either $+1$ or $-1$. Our ablation in the experiments in Sec. VI-C shows that this truncated version alone, which is only sensitive to zeros, works surprisingly well but nevertheless notably worse than the full vector. These experiments also show that the concept hold also across AI codecs, i.e., that JPEG AI can be distinguished from synthetic images even if the features are extracted on channel correlations from the variable bitrate version of the encoder by Ballé et al. [44].
The role of coefficients that are quantized to zero is further illustrated in Table I. Here, we show the average correlation between $\mathbf{y}_{c}$ and $\lfloor\mathbf{y}_{c}\rceil$ for JPEG AI images compressed with various bitrates and synthetic images. For synthetic images we use also the average over Stable Diffusion (XL and V1.4), Glide, Midjourney-V5, Firefly and DALL-E2 images. For stronger JPEG AI compression, the correlation between $\mathbf{y}_{c}$ and $\lfloor\mathbf{y}_{c}\rceil$ decreases. While this is counter-intuitive at first glance, it is a by-product of the quantization. Stronger compression (i.e., lower bpp) leads to more latent space coefficients close to 0, which are rounded to 00 in $\lfloor\mathbf{y}_{c}\rceil$. These are not isolated incidents, but for lower bpps, whole channels $c$ are set to 00, which induces a correlation for that channel of 00. In contrast, synthetic images are much less affected by this effect.
TABLE I: Mean correlation of quantization features for different JPEG AI bitrates (bpps: 0.06 to 0.75) and synthetic images (see text for details). Stronger compression causes lower mean correlation, which seems counter-intuitive, but stems from whole channels being quantized to 00.

### v2

@@ V-A Color Correlations
$\rho(\mathbf{g})=\frac{\langle\delta(\mathbf{r},\mathbf{g}),\delta(\mathbf{g},\mathbf{b})\rangle}{\|\delta(\mathbf{r},\mathbf{g})\|\|\delta(\mathbf{g},\mathbf{b})\|}\kern 5.0pt.$ (4)
@@
Fig. 2 (c) examines correlations induced by traditional JPEG compression. JPEG compression uses a similar preprocessing chain by converting RGB to YCbCr and applying chroma subsampling. The uncompressed patches are compressed with JPEG quality factors 55, 2020, $50$, and $100$. During compression, we ensure that the same 4:2:0 chroma subsampling is used as for JPEG AI. For reasonably high quality factor of $100$, the color correlations slightly increase. With stronger compression of quality factor $50$ the color correlations are almost as much increased as it could be observed for the JPEG AI preprocessing. This is plausible, given the overall similarity of the preprocessing. It is interesting to note that for very strong compressions of quality factors 2020 or even only 55, the correlations are weakened. We did not further investigate into this effect, but we hypothesize that the large quantization factors and associated large rounding errors act as a decorrelator between the color channels.
Fig. 2 (d) examines correlations in synthetic images. Similar correlations in synthetic images have been reported earlier for some color spaces [40]. Here, we show the correlations for synthetic images from the four diffusion-based image generators DALL-E2, Midjourney-V5 (Midj-V5), Firefly, and Stable Diffusion XL (SDXL). Images from DALL-E2 and SDXL exhibit remarkably low correlations in the range of [0.1,+0.15][-0.1,+0.15]. In contrast, images from Midjourney-V5 and Firefly exhibit correlations that are comparable to traditional JPEG compression and JPEG AI compression at very high bitrates. It may be a coincidence that these generators produce images with more natural color appearance, while DALL-E2 and SDXL tend to generate images that oftentimes look oversaturated. In any case, none of the observed correlations from traditional JPEG or synthetic image generators is as large as the correlations for stronger JPEG AI compression.
@@
Figure 4: PSNR between compression k1k-1 and $k$ per bitrate. We analyze JPEG AI (left), the AI codec by Ballé et al. [7] (middle) and JPEG (right). See text for details.
Figure 5: Dependencies between rate $r(k)$ and PSNR $p_{\mathrm{inc}}(k)$ during $k$ compression runs to differentiate between single and recompressed JPEG AI images. In particular, $r(1)$ and the difference r(3)r(2)r(3)-r(2) are more discriminative than the others.
@@ V-B Rate-Distortion Cue
The image and a first version of its associated rate-distortion diagrams are shown in Fig. 3. The left, middle, and right plots show the diagrams for JPEG AI, a further AI codec by Ballé et al. [7], denoted as Ballé et al and traditional JPEG. The $x$-axis indicates the bitrate of the compressed image. JPEG AI allows to directly specify a target bitrate, and for JPEG and Ballé et al. we varied the compression quality to achieve compressions with bitrates between $0.5$ and $3.5$. An image is compressed between 11 (darkblue) to 44 times (lightblue). The $y$-axis indicates the PSNR between the uncompressed input image and the image after $k$ compression steps. For traditional JPEG, repeated compressions barely lead to a change in PSNR, particularly at low bitrates. Only at a bitrate of $2.0$ or higher, PSNRs are decreased by iterated compression. For JPEG AI and Ballé et al., a loss in PSNR upon repeated compression is noticable across the whole range of investigated bitrates.
It is insightful to study a second type of rate-distortion diagram, which is shown in Fig. 4. Here, the $y$-axis indicates the PSNR between two consecutive compressions k1k-1 and $k$. Hence, PSNR is lowest between the original image and the 1st compression, because this compression step changes the image most. Subsequent recompressions have smaller impact, hence the PSNR between compression steps k1k-1 and $k$ is larger. This behavior is analogous for all three codecs, and has also been used in classical works on JPEG recompression forensics [26]. The PSNR differences of both AI codecs increase with increasing bitrates. In contrast, JPEG’s PSNRs are very large at low bitrates, indicating that recompression does not change the image much, but flatten for increasing bitrates somewhere between PSNRs of 45 to 60. As a sidenote, the differences in PSNR in Fig. 4 do not necessarily lead to a noticable difference in PSNR in Fig. 3 (probably best seen for low JPEG bitrates), which is rooted in the instable block phenomenon, where individual JPEG blocks keep changing upon recompression [26, 29].
@@
  • 4 features are obtained from PSNRs. Let $p_{\mathrm{inp}}(k)$ denote the PSNR between the input image and the $k$-th recompression (as in Fig. 3), and let $p_{\mathrm{inc}}(k)$ denote the incremental PSNR between the k1k-1-th and the $k$-th recompression (as in Fig. 4). Then, the features are $p_{\mathrm{inp}}(2)$, $p_{\mathrm{inp}}(3)$, $p_{\mathrm{inc}}(2)$, and $p_{\mathrm{inc}}(3)$.
  • 4 features from the differences (r(3)r(2))(r(3)-r(2)), (r(2)r(1))(r(2)-r(1)), $(p_{\mathrm{inp}}(3)-p_{\mathrm{inp}}(2))$, and $(p_{\mathrm{inc}}(3)-p_{\mathrm{inc}}(2))$.
@@
Figure 5 shows the dependencies between selected features during the compression runs. We observe that the calculated PSNR values and the rate of the three compression runs have a potential to differentiate between single and recompressed JPEG AI images. Notably, $r(1)$ and the difference (r(3)r(2))(r(3)-r(2)) are more discriminative than (r(2)r(1))(r(2)-r(1)) and $(p_{\mathrm{inc}}(3)-p_{\mathrm{inc}}(2))$.
@@ V-C Quantization Cue
These correlations capture two types of properties. First, they are sensitive to zero entries (as a zero entry in one vector cancels the entry in the other vector), and second they tend to reward similar step sizes. Zero entries are arguably most relevant for detecting lossy compression (aiming to maximize sequences of zeros for the arithmetic encoding), but also off-zero entries play an important role. To show this, we also compute a correlation on binarized vectors, i.e., we truncate all non-zero entries to either +1+1 or 1-1. Our ablation in the experiments in Sec. VI-C shows that this truncated version alone, which is only sensitive to zeros, works surprisingly well but nevertheless notably worse than the full vector. These experiments also show that the concept hold also across AI codecs, i.e., that JPEG AI can be distinguished from synthetic images even if the features are extracted on channel correlations from the variable bitrate version of the encoder by Ballé et al. [44].
The role of coefficients that are quantized to zero is further illustrated in Table I. Here, we show the average correlation between $\mathbf{y}_{c}$ and $\lfloor\mathbf{y}_{c}\rceil$ for JPEG AI images compressed with various bitrates and synthetic images. For synthetic images we use also the average over Stable Diffusion (XL and V1.4), Glide, Midjourney-V5, Firefly and DALL-E2 images. For stronger JPEG AI compression, the correlation between $\mathbf{y}_{c}$ and $\lfloor\mathbf{y}_{c}\rceil$ decreases. While this is counter-intuitive at first glance, it is a by-product of the quantization. Stronger compression (i.e., lower bpp) leads to more latent space coefficients close to 0, which are rounded to 0 in $\lfloor\mathbf{y}_{c}\rceil$. These are not isolated incidents, but for lower bpps, whole channels $c$ are set to 0, which induces a correlation for that channel of 0. In contrast, synthetic images are much less affected by this effect.
TABLE I: Mean correlation of quantization features for different JPEG AI bitrates (bpps: 0.06 to 0.75) and synthetic images (see text for details). Stronger compression causes lower mean correlation, which seems counter-intuitive, but stems from whole channels being quantized to 0.